OpenStudy (anonymous):

Can someone please help me with this geometry equation? use the triangle to find the value for cos Ɵ

OpenStudy (anonymous):

here is the picture ^^^

OpenStudy (anonymous):

Are you familiar with SOHCAHTOA?

OpenStudy (zpupster):

do you know the definition for cos

OpenStudy (anonymous):

no, can you remind me?

OpenStudy (zpupster):

sin = opp/hyp cos adj/hyp

OpenStudy (zpupster):

you have the lengths for opp which is 5 and adj is 3 do you know how to find the hypotenuse

OpenStudy (anonymous):

pythagoran therom?

OpenStudy (anonymous):

No, that's why I asked about SOHCAHTOA Sine = Opp/Hyp Cosine = Adj/Hyp Tangent = Opp/Adj They are just special ratios in right triangles

OpenStudy (zpupster):

yes (opp)^2 + (adj)^2 = (hyp)^2 go ahead and plug in you values and solve for hyp then take adj/hyp=cos theta

OpenStudy (anonymous):

@zpupster Why are you telling @Kanine123 to solve for the hypotenuse? You can use the inverse function of the Tangent of Theta to find its value

OpenStudy (zpupster):

yes that is another way

OpenStudy (anonymous):

tan^-1 (theta) = (5/3)

OpenStudy (anonymous):

wait but if you used the Pythagorean theorem and got root 34 for the hypotenuse. would the answer not 3/ root 34?

OpenStudy (anonymous):

@Kanine123 Remember the question is asking for the value of the angle theta, not the hypotenuse, you didn't have to look for the hypotenuse, you don't need that info. in this problem

OpenStudy (zpupster):

use the triangle to find the value for cos Ɵ

OpenStudy (anonymous):

wait but i dont understand, to find the value for cos dont you need the hypotenuse?

OpenStudy (zpupster):

maybe pablo can help you i have ot run

OpenStudy (anonymous):

@Kanine123 This problem does not require any cosines, you need the tangent of theta. This is where SOHCAHTOA comes in handy. TOA, so tangent = opposite over adjacent

OpenStudy (anonymous):

okay so the tangent is 5/3.

OpenStudy (anonymous):

Yeah, so now we know that angle theta has a tangent equal to 5/3 which is roughly 1.6666 repeating. Now we ask ourselves, what angle has that theta ratio? We use a calculator and we type in: Tan-^1( 5/3)

OpenStudy (anonymous):

59.036233 degrees?

OpenStudy (anonymous):

Yeah, that's it. Great! Sometimes we think we have to do extra steps to solve the problem, but that's not necessary xD And remember SOHCAHTOA, it's super handy :D

OpenStudy (anonymous):

awesome! thanks a bunch